OpenStudy (anonymous):

You are designing a new container for powdered laundry detergent. You are considering a cylindrical container with a diameter of 14 inches and a height of 18 inches. Find the volume of this container to the nearest cubic unit. Use a calculator. A. 1,100 in.3 B. 882 in.3 C. 11,084 in.3 D. 2,771 in.3

OpenStudy (anonymous):

help please

OpenStudy (nurali):

Pi r^2 H r=d/2=14/2=7 so 3.14 x 7^2 x 18 = 2769.48 rounded to just 2769

OpenStudy (anonymous):

so is D.2771